How long will a ball take to fall if it starts from rest and moves down for 44.0 m?


1
Expert's answer
2021-09-12T18:59:58-0400

Given:

"h=44.0\:\rm m"


The time of free falling

"t=\sqrt{2h/g}=\sqrt{2*44.0/9.81}=3.00\:\rm s"